• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 00:04:32 on July 23, 2006

    Entry number 174876

    keyword=Shift summary (swing)

    16:00-20:00 Tuning - success at just before 20:00.

    20 microamps of beam provided with what appears to be a better tune. Looked good from our point of view so far: with no target and 20 uA, our HRSs showed no background (trigger rates only 5-10 Hz).

    20:00, MCC asks us to put in our C targets (slanted and slightly-thinner not-slanted) so they can check Ion Chamber rates. It appears that Ion Chamber rates now are about the same as last night, but now we have twice the current as last night.

    22:00-23:00 Did many HARP scans and achieved good beam size and good beam steering.

    23:00 MCC starting to work with Radcon to set Ion Chamber limits to see how high in current we can go.

    23:30 HRS-L Q1 power supply goes off :( Called in Scott to fix - MCC continues working on increasing our current while we wait for Scott. (Thanks Scott!)